The magnitude of the acceleration of an object moving along a straight line is equal to the?

A. slope of the displacement time graph

B. Area below the force-time graph

C. Slope of its velocity-time graph

D. Area under its velocity-time graph


1
Expert's answer
2021-08-18T16:06:22-0400


Option (C) is correct option

Slope of Velocity time graph

Represent of acceleration


The magnitude of the acceleration of an object moving along a straight line is equal to

Slope of velocity time graph




"a=\frac{v_2-v_1}{t_2-t_1}=\frac{∆v}{∆t}"